# Local-global principle for triangularizability and diagonalizability of matrices

### v1

@@ 1. Introduction
  • $\overline{Y_{M,\textup{red}}(k)}=Y_{M,\textup{red}}(\textbf{A}_{k})^{\textup{Br}}_{\bullet}$. In particular, the Brauer–Manin obstruction is the only obstruction to the existence of integral or rational points of $Y_{M,\textup{red}}$.
@@ 2.1. Notation and conventions
For a number field $k$, we denote by $\mathcal{O}_{k}$ the ring of integers of $k$, by $\Omega_{k}$ the set of all places of $k$, and by $\textbf{A}_{k}$ the ring of adèles of $k$. For a finite place $\mathfrak{p}$ of $k$, we denote by $k_{\mathfrak{p}}$ the completion of $k$ at $\mathfrak{p}$, by $\mathrm{ord}_{\mathfrak{p}}(-)$ the valuation associated to $\mathfrak{p}$, and by $\mathcal{O}_{\mathfrak{p}}$ the valuation ring of $k_{\mathfrak{p}}$. For a finite nonempty subset $S$ of $\Omega_{k}$ containing all the archimedean places, we denote by $\mathcal{O}_{k,S}$ the ring of $S$-integers of $k$, which is precisely $\mathcal{O}_{k}$ when $S$ is the set of archimedean places.
@@
Let $V$ be a variety (i.e. a separated scheme of finite type) over a number field $k$, and $\mathcal{V}$ an integral model of $V$ over $\mathcal{O}_{k}$ (i.e. a separated scheme of finite type over $\mathcal{O}_{k}$ whose generic fiber is isomorphic to $V$). We write $V(\textbf{A}_{k})_{\bullet}$ for the reduced adelic space of $V$, where at every real place $v$ the factor $V(k_{v})$ is replaced by its set of connected components, and at every complex place $v$ it is reduced to a single point. The Brauer–Manin set $V(\textbf{A}_{k})^{\textup{Br}}_{\bullet}$, defined as the set of elements of $V(\textbf{A}_{k})_{\bullet}$ that are orthogonal to the Brauer group $\textup{Br}(V)$ of $V$ with respect to the Brauer–Manin pairing, is indeed a closed subset. Moreover, we have a chain of inclusions $V(k)\subset V(\textbf{A}_{k})^{\textup{Br}}_{\bullet}\subset V(\textbf{A}_{k})_{\bullet}$. We say that $V$ satisfies
@@
  • strong approximation if the image of the diagonal map $V(k)\to V(\textbf{A}_{k})_{\bullet}$ is dense.
We say that the Brauer–Manin obstruction is the only obstruction to the existence of rational (resp. integral) points of $V$ if $V(\textbf{A}_{k})^{\textup{Br}}_{\bullet}\neq\varnothing$ implies $V(k)\neq\varnothing$ (resp. $V(\textbf{A}_{k})^{\textup{Br}}_{\bullet}\cap\prod\limits_{v\in\Omega_{k}}\mathcal{V}(\mathcal{O}_{v})\neq\varnothing$ implies $\mathcal{V}(\mathcal{O}_{k})\neq\varnothing$).
@@ 4.1. Diagonalization
  • $\overline{Y_{M,\textup{red}}(k)}=Y_{M,\textup{red}}(\textbf{A}_{k})^{\textup{Br}}_{\bullet}$. In particular, the Brauer–Manin obstruction is the only obstruction to the existence of integral or rational points of $Y_{M,\textup{red}}$.
@@
To show $\overline{Y_{M,\textup{red}}(k)}=Y_{M,\textup{red}}(\textbf{A}_{k})^{\textup{Br}}_{\bullet}$, note that
@@
$Y_{M,\textup{red}}(\textbf{A}_{k})^{\textup{Br}}_{\bullet}=\bigsqcup_{\sigma\in\mathfrak{s}}(Y_{M,\sigma,\textup{red}}(\textbf{A}_{k})^{\textup{Br}}_{\bullet})=\bigsqcup_{\sigma\in\mathfrak{s}}(\mathrm{GL}_{n_{1}}(\textbf{A}_{k})^{\textup{Br}}_{\bullet}\times...\times\mathrm{GL}_{n_{r}}(\textbf{A}_{k})^{\textup{Br}}_{\bullet})$
@@ 4.2. Triangularization
Based on previous propositions, we know that the arithmetic properties of $X_{M,\textup{red}}$ reduce to $X_{J(\lambda_{i}),\textup{red}}$. In particular, if $\overline{X_{J(\lambda_{i}),\textup{red}}(k)}=X_{J(\lambda_{i}),\textup{red}}(\textbf{A}_{k})^{\textup{Br}}_{\bullet}$ for each $1\leq i\leq r$ then $\overline{X_{M,\textup{red}}(k)}=X_{M,\textup{red}}(\textbf{A}_{k})^{\textup{Br}}_{\bullet}$, hence the Brauer–Manin obstruction will be the only obstruction to the existence of integral or rational points of $X_{M,\textup{red}}$ by the same proof of the case of diagonalization.
@@
In both cases below, we have $\overline{X_{J(\lambda_{i}),\textup{red}}(k)}=X_{J(\lambda_{i}),\textup{red}}(\textbf{A}_{k})^{\textup{Br}}_{\bullet}$.
@@ 5. The stratified Brauer–Manin obstruction
The stratified Brauer–Manin set of $V$ is defined as $\bigcup_{i}V_{i}(\textbf{A}_{k})_{\bullet}^{\textup{Br}}$, and we denote it by $V(\textbf{A}_{k})_{\bullet}^{\mathrm{st}\mathrm{Br}}$. Note that the following inclusions: $V(k)\subset V(\textbf{A}_{k})_{\bullet}^{\mathrm{st}\mathrm{Br}}\subset V_{\textup{red}}(\textbf{A}_{k})_{\bullet}^{\textup{Br}}\subset V(\textbf{A}_{k})_{\bullet}$.
We shall say that the stratified Brauer–Manin obstruction is the only obstruction to the existence of rational (resp. integral) points of $V$ if $V(\textbf{A}_{k})_{\bullet}^{\mathrm{st}\mathrm{Br}}\neq\varnothing$ implies $V(k)\neq\varnothing$ (resp. $V(\textbf{A}_{k})_{\bullet}^{\mathrm{st}\mathrm{Br}}\bigcap\prod\limits_{v\in\Omega_{k}}\mathcal{V}(\mathcal{O}_{v})\neq\varnothing$ implies $\mathcal{V}(\mathcal{O}_{k})\neq\varnothing$).
@@
If $M\in\mathrm{M}_{m+n}(\mathcal{O}_{k})$ is similar to $\begin{pmatrix}\lambda I_{m}&0\\ 0&J_{n}(\lambda)\end{pmatrix}$ over $k$, where $\lambda\in\mathcal{O}_{k}$ and $J_{n}(\lambda)$ is the $n\times n$ Jordan block with eigenvalue $\lambda$, then Conjecture 5.2 and 5.3 are true.
The rest of this section is devoted to the proof of Theorem 5.4. Without loss of generality, we may assume that $M=\mathrm{diag}(\lambda I_{m},J_{n}(\lambda))$, since similarity of matrices induces an isomorphism between the corresponding varieties. The elementary cases $M=\lambda I_{m}$ and $M=J_{n}(\lambda)$ have already been settled in Proposition 4.5. In what follows, we therefore restrict to the situation where $m\geq 1$ and $n\geq 2$.
@@
  • for every $x\in V(k)$, the fiber $W_{x}$ satisfies $\overline{W_{x}(k)}=W_{x}(\textbf{A}_{k})_{\bullet}$.
In fact, the second condition is satisfied when all fibers of $f$ are nonempty and geometrically connected, cf. [Con12, Theorem 4.5]. In what follows, unless otherwise specified, we shall always use the fact that all fibers are geometrically integral to verify the second condition stated above.
@@
    $\overline{V(k)}=V(\textbf{A}_{k})_{\bullet}^{\textup{Br}}$ implies $\overline{W(k)}=W(\textbf{A}_{k})_{\bullet}^{\textup{Br}}$,
@@
    $\overline{V(k)}=V(\textbf{A}_{k})_{\bullet}$ implies $\overline{W(k)}=W(\textbf{A}_{k})_{\bullet}$.
@@
    Without loss of generality, we assume that $W(k)$ is nonempty. Then take a nonempty open subset $\Delta$ of $W(\textbf{A}_{k})_{\bullet}^{\textup{Br}}$, we want to show that $W(k)\cap\Delta\neq\varnothing$. We may assume $\Delta$ is of the form $(\prod\limits_{v\in T}U_{v}\times\prod\limits_{v\notin T}\mathcal{W}(\mathcal{O}_{v}))\cap W(\textbf{A}_{k})_{\bullet}^{\textup{Br}}$, where $T$ is a finite set of places containing $S$, $U_{v}$ is an open subset of $W(k_{v})$. Also there exists a point $(\alpha_{v})_{v\in T}\times(\beta_{v})_{v\notin T}\in\prod\limits_{v\in T}U_{v}\times\prod\limits_{v\notin T}\mathcal{W}(\mathcal{O}_{k})$ that is orthogonal to $\textup{Br}(W)$.
    By implicit function theorem, $f(U_{v})\subset V(k_{v})$ is an open subset, then $(\prod\limits_{v\in T}f(U_{v})\times\prod\limits_{v\notin T}\mathcal{V}(\mathcal{O}_{v}))\cap V(\textbf{A}_{k})_{\bullet}^{\textup{Br}}$ is an open subset of $V(\textbf{A}_{k})_{\bullet}^{\textup{Br}}$ containing $(f(\alpha_{v}))_{v\in T}\times(f(\beta_{v}))_{v\notin T}$. Hence there exists an element $x\in V(k)$ contained in $\prod\limits_{v\in T}f(U_{v})\times\prod\limits_{v\notin T}\mathcal{V}(\mathcal{O}_{v})$. The fiber $W_{x}$ satisfies $\overline{W_{x}(k)}=W_{x}(\textbf{A}_{k})_{\bullet}$. Moreover $x$ is an $\mathcal{O}_{v}$-point of $\mathcal{V}$, we can define $\mathcal{W}_{x,v}$ as the fiber of $\mathcal{W}$ over this $\mathcal{O}_{v}$-point. By construction we know that $\prod\limits_{v\in T}(U_{v}\cap W_{x}(k_{v}))\times\prod\limits_{v\notin T}\mathcal{W}_{x,v}(\mathcal{O}_{v})$ is a nonempty open subset of $W_{x}(\textbf{A}_{k})_{\bullet}$. Therefore there exists an element $x^{\prime}\in W_{x}(k)$ contained in $\prod\limits_{v\in T}(U_{v}\cap W_{x}(k_{v}))\times\prod\limits_{v\notin T}\mathcal{W}_{x,v}(\mathcal{O}_{v})\subset\Delta$. This completes the proof.
@@
    $W=\textup{Spec}\ k[T_{i,j},1\leq i\leq n,1\leq j\leq m_{i}]/(1-\sum\limits_{1\leq i\leq n}\prod\limits_{1\leq j\leq m_{i}}T_{i,j})$ satisfies $\overline{W(k)}=W(\textbf{A}_{k})_{\bullet}$,
@@
    $W^{\prime}=\textup{Spec}\ k[T_{i,j},1\leq i\leq n,1\leq j\leq m_{i};\frac{1}{\sum\limits_{1\leq i\leq n}\prod\limits_{1\leq j\leq m_{i}}T_{i,j}}]$ satisfies $\overline{W^{\prime}(k)}=W^{\prime}(\textbf{A}_{k})_{\bullet}^{\textup{Br}}$.
@@
$(1)$ Consider the morphism $f:W\to V=\textup{Spec}\ k[Z_{i,j},1\leq i\leq n,2\leq j\leq m_{i}]$-$\square$, which maps $Z_{i,j}$ to $T_{i,j}$, where $\square\subset\textup{Spec}\ k[Z_{i,j},1\leq i\leq n,2\leq j\leq m_{i}]$ is the closed subscheme defined by $(\prod\limits_{2\leq j\leq m_{i}}T_{1,j})=...=(\prod\limits_{2\leq j\leq m_{i}}T_{n,j})=0$. Then $\overline{V(k)}=V(\textbf{A}_{k})_{\bullet}$ by Lemma 2.3. One can verify that $f$ is a good fibration. Indeed, the fibre of $f$ over any $k$-point has constant dimension; that is, all fibres are isomorphic to affine spaces of the same dimension. By miracle flatness, it follows that $f$ is flat, and hence smooth. Moreover, the surjectivity on integral points arises naturally from the construction.
@@
For each integer $n\geq 1$, and $i,j$ such that $1\leq j\leq i\leq n$, we define a set $\mathfrak{M}_{n}(i,j)$ as a subset of $k[T_{i,j}\mid_{1\leq i,j\leq n}]$ which contains all minors of order $j$ obtained by choosing any $j$ rows among the first $i$ rows and taking the first $j$ columns of the matrix $\left(T_{i,j}\right)_{n\times n}$.
@@
For positive integers $r,s$, such that $r<s$, we have $\overline{\mathrm{SL}_{s,r}(k)}=\mathrm{SL}_{s,r}(\textbf{A}_{k})_{\bullet}$.
@@
One checks that $f$ is a good fibration whose fibers are always isomorphic to $\mathbb{A}^{s-r}$. Moreover, $r<s$ implies that $\cap_{r\leq l\leq s}V(C_{s,l})\subset\textup{Spec}\ k[w_{i,j}\mid 1\leq i\leq s-1,1\leq j\leq s]$ is a closed subscheme of codimension at least $2$. Therefore, we have $\overline{\mathrm{SL}_{s,r}(k)}=\mathrm{SL}_{s,r}(\textbf{A}_{k})_{\bullet}$ by Lemma 2.3 and Lemma 5.6. ∎
@@
Then $W_{\mathbf{s}}$ is a smooth and $k$-rational variety satisfying $\overline{W_{\mathbf{s}}(k)}=W_{\mathbf{s}}(\textbf{A}_{k})_{\bullet}$.
Similarly, let $W^{\prime}_{\mathbf{s}}$ be the closed subscheme of $\textup{Spec}\ k[T_{i,j}\mid_{1\leq i,j\leq m+n};\frac{1}{\det(T_{i,j})}]$ defined by the same equations as above, then $W^{\prime}_{\mathbf{s}}$ is a smooth and $k$-rational variety satisfying $\overline{W^{\prime}_{\mathbf{s}}(k)}=W^{\prime}_{\mathbf{s}}(\textbf{A}_{k})_{\bullet}^{\textup{Br}}$.
@@
$f:W_{\textbf{s}}\to\textup{Spec}\ k[X_{s_{n-2}+1},...,X_{s_{n-1}};Y_{s_{n-2}+1},...,Y_{s_{n-1}};\frac{1}{X_{s_{n-2}+1}Y_{s_{n-2}+1}+...+X_{s_{n-1}}Y_{s_{n-1}}}],$
@@
are linearly independent. Hence “$f=0$ for any $f\in\mathfrak{M}_{m+n}(m+n-2,s_{n-1})$” implies that $\begin{pmatrix}T_{1,\beta}&\cdots&T_{m+n-2,\beta}\end{pmatrix}^{\mathrm{T}}$ is a linear combination of the above vectors. Therefore, when $n=2$, the fiber of $f$ at every $k$-point is isomorphic to $\mathrm{SL}_{m+1,s_{2}-1}\times\mathbb{A}^{m+2-s_{1}}$, hence $f$ is a good fibration, and $\overline{W_{\textbf{s}}(k)}=W_{\textbf{s}}(\textbf{A}_{k})_{\bullet}$ follows from Corollary 5.7.
Similarly, when $n\geq 3$, the fiber of $f$ at every $k$-point is isomorphic to $W^{\prime}_{\textbf{s}^{\prime}}\times\mathbb{A}^{m+n-2-s_{n-1}}$, where $\textbf{s}^{\prime}=(s_{1},...,s_{n-2},s_{n}-1)\in\mathfrak{S}_{m,n-1}$ is a well-defined element. It then follows from the induction hypothesis that $f$ is a good fibration and $\overline{W_{\textbf{s}}(k)}=W_{\textbf{s}}(\textbf{A}_{k})_{\bullet}$. This completes the proof. ∎

### v2

@@ 1. Introduction
  • $\overline{Y_{M,\textup{red}}(k)}=Y_{M,\textup{red}}(\mathbf{A}_{k})^{\textup{Br}}_{\bullet}$. In particular, the Brauer–Manin obstruction is the only obstruction to the existence of integral or rational points of $Y_{M,\textup{red}}$.
@@ 2.1. Notation and conventions
For a number field $k$, we denote by $\mathcal{O}_{k}$ the ring of integers of $k$, by $\Omega_{k}$ the set of all places of $k$, and by $\mathbf{A}_{k}$ the ring of adèles of $k$. For a finite place $\mathfrak{p}$ of $k$, we denote by $k_{\mathfrak{p}}$ the completion of $k$ at $\mathfrak{p}$, by $\mathrm{ord}_{\mathfrak{p}}(-)$ the valuation associated to $\mathfrak{p}$, and by $\mathcal{O}_{\mathfrak{p}}$ the valuation ring of $k_{\mathfrak{p}}$. For a finite nonempty subset $S$ of $\Omega_{k}$ containing all the archimedean places, we denote by $\mathcal{O}_{k,S}$ the ring of $S$-integers of $k$, which is precisely $\mathcal{O}_{k}$ when $S$ is the set of archimedean places.
@@
Let $V$ be a variety (i.e. a separated scheme of finite type) over a number field $k$, and $\mathcal{V}$ an integral model of $V$ over $\mathcal{O}_{k}$ (i.e. a separated scheme of finite type over $\mathcal{O}_{k}$ whose generic fiber is isomorphic to $V$). We write $V(\mathbf{A}_{k})_{\bullet}$ for the reduced adelic space of $V$, where at every real place $v$ the factor $V(k_{v})$ is replaced by its set of connected components, and at every complex place $v$ it is reduced to a single point. The Brauer–Manin set $V(\mathbf{A}_{k})^{\textup{Br}}_{\bullet}$, defined as the set of elements of $V(\mathbf{A}_{k})_{\bullet}$ that are orthogonal to the Brauer group $\textup{Br}(V)$ of $V$ with respect to the Brauer–Manin pairing, is indeed a closed subset. Moreover, we have a chain of inclusions $V(k)\subset V(\mathbf{A}_{k})^{\textup{Br}}_{\bullet}\subset V(\mathbf{A}_{k})_{\bullet}$. We say that $V$ satisfies
@@
  • strong approximation if the image of the diagonal map $V(k)\to V(\mathbf{A}_{k})_{\bullet}$ is dense.
We say that the Brauer–Manin obstruction is the only obstruction to the existence of rational (resp. integral) points of $V$ if $V(\mathbf{A}_{k})^{\textup{Br}}_{\bullet}\neq\varnothing$ implies $V(k)\neq\varnothing$ (resp. $V(\mathbf{A}_{k})^{\textup{Br}}_{\bullet}\cap\prod\limits_{v\in\Omega_{k}}\mathcal{V}(\mathcal{O}_{v})\neq\varnothing$ implies $\mathcal{V}(\mathcal{O}_{k})\neq\varnothing$).
@@ 4.1. Diagonalization
  • $\overline{Y_{M,\textup{red}}(k)}=Y_{M,\textup{red}}(\mathbf{A}_{k})^{\textup{Br}}_{\bullet}$. In particular, the Brauer–Manin obstruction is the only obstruction to the existence of integral or rational points of $Y_{M,\textup{red}}$.
@@
To show $\overline{Y_{M,\textup{red}}(k)}=Y_{M,\textup{red}}(\mathbf{A}_{k})^{\textup{Br}}_{\bullet}$, note that
@@
$Y_{M,\textup{red}}(\mathbf{A}_{k})^{\textup{Br}}_{\bullet}=\bigsqcup_{\sigma\in\mathfrak{s}}(Y_{M,\sigma,\textup{red}}(\mathbf{A}_{k})^{\textup{Br}}_{\bullet})=\bigsqcup_{\sigma\in\mathfrak{s}}(\mathrm{GL}_{n_{1}}(\mathbf{A}_{k})^{\textup{Br}}_{\bullet}\times...\times\mathrm{GL}_{n_{r}}(\mathbf{A}_{k})^{\textup{Br}}_{\bullet})$
@@ 4.2. Triangularization
Based on previous propositions, we know that the arithmetic properties of $X_{M,\textup{red}}$ reduce to $X_{J(\lambda_{i}),\textup{red}}$. In particular, if $\overline{X_{J(\lambda_{i}),\textup{red}}(k)}=X_{J(\lambda_{i}),\textup{red}}(\mathbf{A}_{k})^{\textup{Br}}_{\bullet}$ for each $1\leq i\leq r$ then $\overline{X_{M,\textup{red}}(k)}=X_{M,\textup{red}}(\mathbf{A}_{k})^{\textup{Br}}_{\bullet}$, hence the Brauer–Manin obstruction will be the only obstruction to the existence of integral or rational points of $X_{M,\textup{red}}$ by the same proof of the case of diagonalization.
@@
In both cases below, we have $\overline{X_{J(\lambda_{i}),\textup{red}}(k)}=X_{J(\lambda_{i}),\textup{red}}(\mathbf{A}_{k})^{\textup{Br}}_{\bullet}$.
@@ 5. The stratified Brauer–Manin obstruction
The stratified Brauer–Manin set of $V$ is defined as $\bigcup_{i}V_{i}(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}$, and we denote it by $V(\mathbf{A}_{k})_{\bullet}^{\mathrm{st}\mathrm{Br}}$. Note that the following inclusions: $V(k)\subset V(\mathbf{A}_{k})_{\bullet}^{\mathrm{st}\mathrm{Br}}\subset V_{\textup{red}}(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}\subset V(\mathbf{A}_{k})_{\bullet}$.
We shall say that the stratified Brauer–Manin obstruction is the only obstruction to the existence of rational (resp. integral) points of $V$ if $V(\mathbf{A}_{k})_{\bullet}^{\mathrm{st}\mathrm{Br}}\neq\varnothing$ implies $V(k)\neq\varnothing$ (resp. $V(\mathbf{A}_{k})_{\bullet}^{\mathrm{st}\mathrm{Br}}\bigcap\prod\limits_{v\in\Omega_{k}}\mathcal{V}(\mathcal{O}_{v})\neq\varnothing$ implies $\mathcal{V}(\mathcal{O}_{k})\neq\varnothing$).
@@
If $M\in\mathrm{M}_{m+n}(\mathcal{O}_{k})$ is similar to $\begin{pmatrix}\lambda I_{m}&0\\ 0&J_{n}(\lambda)\end{pmatrix}$ over $k$ for some $\lambda\in\mathcal{O}_{k}$, then Conjecture 5.2 and Conjecture 5.3 are true.
The rest of this section is devoted to the proof of Theorem 5.4. Without loss of generality, we may assume that $M=\mathrm{diag}(\lambda I_{m},J_{n}(\lambda))$, since similar matrices correspond to isomorphic varieties. The elementary cases $M=\lambda I_{m}$ and $M=J_{n}(\lambda)$ have already been settled in Proposition 4.5. In what follows, we therefore restrict to the situation where $m\geq 1$ and $n\geq 2$.
@@
  • for every $x\in V(k)$, the fiber $W_{x}$ satisfies $\overline{W_{x}(k)}=W_{x}(\mathbf{A}_{k})_{\bullet}$.
In fact, the second condition is satisfied when all fibers of $f$ are nonempty and geometrically connected, cf. [Con12, Theorem 4.5]. In what follows, we shall always use the fact that all fibers are geometrically integral to verify the second condition stated above.
@@
    $\overline{V(k)}=V(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}$ implies $\overline{W(k)}=W(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}$,
@@
    $\overline{V(k)}=V(\mathbf{A}_{k})_{\bullet}$ implies $\overline{W(k)}=W(\mathbf{A}_{k})_{\bullet}$.
@@
    Without loss of generality, we assume that $W(k)$ is nonempty. Then take a nonempty open subset $\Delta$ of $W(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}$, we want to show that $W(k)\cap\Delta\neq\varnothing$. We may assume $\Delta$ is of the form $(\prod\limits_{v\in T}U_{v}\times\prod\limits_{v\notin T}\mathcal{W}(\mathcal{O}_{v}))\cap W(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}$, where $T$ is a finite set of places containing $S$, $U_{v}$ is an open subset of $W(k_{v})$. Also there exists a point $(\alpha_{v})_{v\in T}\times(\beta_{v})_{v\notin T}\in\prod\limits_{v\in T}U_{v}\times\prod\limits_{v\notin T}\mathcal{W}(\mathcal{O}_{k})$ that is orthogonal to $\textup{Br}(W)$.
    By implicit function theorem, $f(U_{v})\subset V(k_{v})$ is an open subset, then $(\prod\limits_{v\in T}f(U_{v})\times\prod\limits_{v\notin T}\mathcal{V}(\mathcal{O}_{v}))\cap V(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}$ is an open subset of $V(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}$ containing $(f(\alpha_{v}))_{v\in T}\times(f(\beta_{v}))_{v\notin T}$. Hence there exists an element $x\in V(k)$ contained in $\prod\limits_{v\in T}f(U_{v})\times\prod\limits_{v\notin T}\mathcal{V}(\mathcal{O}_{v})$. The fiber $W_{x}$ satisfies $\overline{W_{x}(k)}=W_{x}(\mathbf{A}_{k})_{\bullet}$. Moreover $x$ is an $\mathcal{O}_{v}$-point of $\mathcal{V}$, we can define $\mathcal{W}_{x,v}$ as the fiber of $\mathcal{W}$ over this $\mathcal{O}_{v}$-point. By construction we know that $\prod\limits_{v\in T}(U_{v}\cap W_{x}(k_{v}))\times\prod\limits_{v\notin T}\mathcal{W}_{x,v}(\mathcal{O}_{v})$ is a nonempty open subset of $W_{x}(\mathbf{A}_{k})_{\bullet}$. Therefore there exists an element $x^{\prime}\in W_{x}(k)$ contained in $\prod\limits_{v\in T}(U_{v}\cap W_{x}(k_{v}))\times\prod\limits_{v\notin T}\mathcal{W}_{x,v}(\mathcal{O}_{v})\subset\Delta$. This completes the proof.
@@
    $W=\textup{Spec}\ k[T_{i,j},1\leq i\leq n,1\leq j\leq m_{i}]/(1-\sum\limits_{1\leq i\leq n}\prod\limits_{1\leq j\leq m_{i}}T_{i,j})$ satisfies $\overline{W(k)}=W(\mathbf{A}_{k})_{\bullet}$,
@@
    $W^{\prime}=\textup{Spec}\ k[T_{i,j},1\leq i\leq n,1\leq j\leq m_{i};\frac{1}{\sum\limits_{1\leq i\leq n}\prod\limits_{1\leq j\leq m_{i}}T_{i,j}}]$ satisfies $\overline{W^{\prime}(k)}=W^{\prime}(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}$.
@@
$(1)$ Consider the morphism $f:W\to V=\textup{Spec}\ k[Z_{i,j},1\leq i\leq n,2\leq j\leq m_{i}]$-$\square$, which maps $Z_{i,j}$ to $T_{i,j}$, where $\square\subset\textup{Spec}\ k[Z_{i,j},1\leq i\leq n,2\leq j\leq m_{i}]$ is the closed subscheme defined by $(\prod\limits_{2\leq j\leq m_{i}}T_{1,j})=...=(\prod\limits_{2\leq j\leq m_{i}}T_{n,j})=0$. Then $\overline{V(k)}=V(\mathbf{A}_{k})_{\bullet}$ by Lemma 2.3. One can verify that $f$ is a good fibration. Indeed, the fibre of $f$ over any $k$-point has constant dimension; that is, all fibres are isomorphic to affine spaces of the same dimension. By miracle flatness, it follows that $f$ is flat, and hence smooth. Moreover, the surjectivity on locally integral points arises from the geometrically integral fibres.
@@
For each integer $n\geq 1$, and $r,s$ such that $1\leq r\leq s\leq n$, we define a set $\mathfrak{M}_{n}(s,r)$ as a subset of $k[T_{i,j}\mid_{1\leq i,j\leq n}]$ which contains all minors of order $r$ obtained by choosing any $r$ rows among the first $s$ rows and taking the first $r$ columns of the matrix $\left(T_{i,j}\right)_{n\times n}$.
@@
For positive integers $r,s$, such that $r<s$, we have $\overline{\mathrm{SL}_{s,r}(k)}=\mathrm{SL}_{s,r}(\mathbf{A}_{k})_{\bullet}$.
@@
One checks that $f$ is a good fibration whose fibers are always isomorphic to $\mathbb{A}^{s-r}$. Moreover, $r<s$ implies that $\cap_{r\leq l\leq s}V(C_{s,l})\subset\textup{Spec}\ k[w_{i,j}\mid 1\leq i\leq s-1,1\leq j\leq s]$ is a closed subscheme of codimension at least $2$. Therefore, we have $\overline{\mathrm{SL}_{s,r}(k)}=\mathrm{SL}_{s,r}(\mathbf{A}_{k})_{\bullet}$ by Lemma 2.3 and Lemma 5.6. ∎
@@
Then $W_{\mathbf{s}}$ is a smooth and $k$-rational variety satisfying $\overline{W_{\mathbf{s}}(k)}=W_{\mathbf{s}}(\mathbf{A}_{k})_{\bullet}$.
Similarly, let $W^{\prime}_{\mathbf{s}}$ be the closed subscheme of $\textup{Spec}\ k[T_{i,j}\mid_{1\leq i,j\leq m+n};\frac{1}{\det(T_{i,j})}]$ defined by the same equations as above, then $W^{\prime}_{\mathbf{s}}$ is a smooth and $k$-rational variety satisfying $\overline{W^{\prime}_{\mathbf{s}}(k)}=W^{\prime}_{\mathbf{s}}(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}$.
@@
$f:W_{\textbf{s}}\to\textup{Spec}\ k[X_{s_{n-2}+1},...,X_{s_{n-1}};Y_{s_{n-2}+1},...,Y_{s_{n-1}}]/(1-\sum_{i=s_{n-2}+1}^{s_{n-1}}X_{i}Y_{i}),$
@@
are linearly independent. Hence “$f=0$ for any $f\in\mathfrak{M}_{m+n}(m+n-2,s_{n-1})$” implies that $\begin{pmatrix}T_{1,\beta}&\cdots&T_{m+n-2,\beta}\end{pmatrix}^{\mathrm{T}}$ is a linear combination of the above vectors. Therefore, when $n=2$, each fiber of $f$ is isomorphic to $\mathrm{SL}_{m+1,s_{2}-1}\times\mathbb{A}^{m+2-s_{1}}$, hence $f$ is a good fibration, and $\overline{W_{\textbf{s}}(k)}=W_{\textbf{s}}(\mathbf{A}_{k})_{\bullet}$ follows from Corollary 5.7.
Similarly, when $n\geq 3$, each fiber of $f$ is isomorphic to $W_{\textbf{s}^{\prime}}\times\mathbb{A}^{m+n-s_{n-1}}$, where $\textbf{s}^{\prime}=(s_{1},...,s_{n-2},s_{n}-1)\in\mathfrak{S}_{m,n-1}$ is a well-defined element. It then follows from the induction hypothesis that $f$ is a good fibration and $\overline{W_{\textbf{s}}(k)}=W_{\textbf{s}}(\mathbf{A}_{k})_{\bullet}$. This completes the proof. ∎
@@
The proof of Lemma 5.14 is somewhat technical and is deferred to Appendix A for the sake of readability.
Lemma 5.15.
[…]
Then each $V_{\textbf{r}}$ is a closed subscheme of $X_{M}$. Moreover, we have
$X_{M}=\bigcup_{\textbf{r}\in\mathfrak{R}_{m,n}}V_{\textbf{r}}.$